OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4. y = x + 2 y = -x - 2 y = -x + 4 y = x - 4

OpenStudy (anonymous):

let x=-3 then y=? 1) y=-3+2=-1 2) y=-(-3)-2=3-2=1 3) y=-(-3)+4=3+4=7 4) y=-3-4=-7

OpenStudy (lena772):

B